OpenStudy (kingstone):

Find the specified vector or scalar. u = -4i + 1j and v = 4i + 1j ; Find ||u+v||. A. square root of 34. B. 8 C. 2 D. 5

OpenStudy (anonymous):

2

OpenStudy (anonymous):

need explanation?

OpenStudy (kingstone):

u+v = 2j and ||u+v|| = \[\sqrt{a ^{2} + b^{2}}\] so ||u+v|| = square root of 2^2 = 2 I got the answer myself already, but your approval made me sure of my answer, so Thank You very much!!!

OpenStudy (anonymous):

good luck :)